# Global Behavior of Learning Dynamics in Zero-Sum Games with Memory Asymmetry

### v1

@@ 3.4 Family of Lyapunov functions
This means that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ for all $\bm{\delta}$. ∎
Let us interpret the function of $H(\bm{X};\bm{\delta})$. For simplicity, we consider a simple class of zero-sum vectors, $\bm{\delta}=\bm{e}_{i}-\bm{e}_{i^{\prime}}$ for some $i$ and $i^{\prime}\neq i$, where $\bm{e}_{i}$ denotes the unit vector for $i$-th element. Then, we calculate the function as $H(\bm{X};\bm{e}_{i}-\bm{e}_{i^{\prime}})=\bm{u}_{i}\cdot\log\bm{x}_{i}-\bm{u}_{i}\cdot\log\bm{x}_{i^{\prime}}-\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i}+\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i^{\prime}}$. The monotonic increase of $H(\bm{X};\bm{e}_{i}-\bm{e}_{i^{\prime}})$ shows that $\log\bm{x}_{i}$ becomes correlated to $\bm{u}_{i}$, and $\log\bm{x}_{i^{\prime}}$ to $\bm{u}_{i^{\prime}}$. Here, $\bm{x}_{i}=(x_{i|j})_{j}$ represents the vector of the probabilities to choose $i$-th action in the condition that the other chose $j$-th action in the previous round. $\bm{u}_{i}=(u_{ij})_{j}$ also represents the vector of the payoffs for action $i$. Thus, the correlation between $\log\bm{x}_{i}$ and $\bm{u}_{i}$ means that the larger its reward $u_{ij}$ is, the more frequently X returns its $i$-th action against the other’s $j$-th action, in other words, X becomes more exploitative.
This theorem characterizes the state of learning dynamics after a sufficiently long time passes, as follows.
Corollary 2 (Convergence of Y’s strategy or divergence of $H$).
After sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds or $H(\bm{X};\bm{\delta})$ diverges to infinity for all $\bm{\delta}\neq\bm{0}$ such that $\Sigma_{i}\delta_{i}=0$.
Proof.
Since $H(\bm{X};\bm{\delta})\geq 0$ always holds for all $\bm{\delta}$, either $H(\bm{X};\bm{\delta})=0$ or $H(\bm{X};\bm{\delta})>0$ holds in the final state. The former is equivalent to $\bm{y}=\bm{y}^{*}$, while the latter is equivalent to $H(\bm{X};\bm{\delta})\to\infty$. ∎
4 Global behavior of learning dynamics
Now, Cors. 1 and 2 are utilized to capture the global behavior of the learning dynamics. One of the global behaviors seen in the zero-sum games is the global convergence to the Nash equilibrium independent of X’s and Y’s initial strategies. First, Cor. 1 guarantees that after a sufficiently long time passes, either Y’s strategy converges to the equilibrium or $H(\bm{X};\bm{\delta})$ diverges. Since the former trivially leads to such global convergence, we now consider the latter condition, i.e., $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}\to\infty$. Here, $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}\to\infty$ naively connects to $-D^{\dagger}(\bm{X},{\rm d}\bm{y})={\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}>0$ (explained in the next paragraph), meaning that $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors. If so, Cor. 2 becomes applicable, and either Y’s strategy converges to the Nash equilibrium.
We explain the connection between $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}$ and $-D^{\dagger}(\bm{X},{\rm d}\bm{y})={\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}>0$. Here, $H(\bm{X};\bm{\delta})$ and $-D^{\dagger}(\bm{X},{\rm d}\bm{y})$ have two commonalities. The first one is that $\bm{\delta}$ and ${\rm d}\bm{y}$ are both zero-sum vectors in common. The other one is that both $H$ and $-D^{\dagger}$ are given by the product of X’s strategy $\bm{X}$ and the payoff matrix $\bm{U}$. This means that the more $\bm{X}$ is correlated to $\bm{U}$, the larger both $H$ and $-D^{\dagger}$ are. Here, however, we also remark two differences of $H$ from $-D^{\dagger}$; $H$ considers the logarithm in X’s strategy $\bm{X}$ and the reversed order in the product of $\bm{X}$ and $\bm{U}$.
In the following, we prove such global convergence of either player’s strategy in the matching pennies, the simplest game equipped with a full-support equilibrium. We also experimentally show global convergence in more complicated games than the matching pennies later.
4.1 Example: Matching-pennies
Let us define the matching-pennies game (see Fig. 2 for the illustration of its payoff matrix). This game considers the action numbers of $m_{\sf X}=m_{\sf Y}=2$ and the payoff matrix of $\bm{U}=((u_{11},u_{12}),(u_{21},u_{22}))=((+1,-1),(-1,+1))$. The Nash equilibrium of this game is only $\bm{x}^{*}=\bm{y}^{*}=(1/2,1/2)$. This game is the simplest example of games that have a full-support equilibrium. In addition, it has been known that the replicator dynamics in games without memories cycle around the Nash equilibrium and cannot reach the equilibrium. Nevertheless, the learning dynamics in games of a reactive strategy achieve the convergence to the Nash equilibrium for Y’s strategy. The following theorem holds (see Appendix A.3 for its full proof).
Theorem 3 (Global convergence in matching-pennies).
In the matching-pennies ($\bm{U}=((+1,-1),(-1,+1))$), Y’s strategy converges to the Nash equilibrium ($\bm{y}=\bm{y}^{*}$), independent of both the players’ initial strategies.
Proof Sketch. First, we define a special notation in two-action games; $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$ for $j=1,2$. We also define the logit function of $x_{j}$ as $q_{j}:=\log x_{j}-\log(1-x_{j})$. Here, note that $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$. By the direct calculation, we prove $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$. Thus, Cor. 2 shows that after a sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds or $H(\bm{X};\bm{\delta})$ diverges, in other words, $H(\bm{X};\bm{\delta})>0\Leftrightarrow q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ holds. In the former case, Y obviously reaches the Nash equilibrium strategy, so we now assume the latter case ($x_{1}>x_{2}$ holds). We can also prove that $x_{1}>x_{2}$ is equivalent to the positive definiteness of $\bm{X}^{\rm T}\bm{U}$. Because $\bm{X}^{\rm T}\bm{U}$ continues to hold after a sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds after a further long time passes, meaning that Y’s strategy converges to its equilibrium. ∎
5 Experimental supports for theory
This section experimentally shows the global convergence to the Nash equilibrium. First, we visualize the mechanism of global convergence in matching pennies. Following Thm. 2, $H(\bm{X};\bm{\delta})$ increases with time, and $\bm{X}^{\rm T}\bm{U}$ becomes positive definite. After that, following Thm. 1, the distance from the Nash equilibrium $D(\bm{X},\bm{y})$ decreases with time. Next, we also demonstrate global convergence for several examples of more advanced games than matching pennies. We have observed the global convergence in all zero-sum games as far as we experimented, even though we could not prove it in general zero-sum games. All the simulations in this section are performed following the Runge-Kutta fourth-order method of Eqs. (5) and (6) with a sufficiently small step size of $5\times 10^{-2}$.
@@ 5 Experimental supports for theory
Figure 3: A. Trajectories of $q_{1}$ and $q_{2}$. The rainbow contour plot indicates the value of $q_{1}-q_{2}$. All the trajectories monotonically increase $q_{1}-q_{2}$ with time and converge in the area of $q_{1}>q_{2}$ in their final states. B. Trajectories of the learning dynamics. The black broken line corresponds to the region of Nash equilibria, $\bm{x}^{{\rm st}}=\bm{y}=(1/2,1/2)$. Each colored line shows a trajectory of the learning dynamics. First, the circle markers show the initial states. Following the blue lines, the trajectories diverge from the Nash equilibria ($D(\bm{X},\bm{y})$ increases with time). However, the trajectories stop to diverge and switch to converge to the Nash equilibria ($D(\bm{X},\bm{y})$ decreases), following the red lines. The star markers are the final states and correspond to one of the Nash equilibria.
5.1 Visualization of global convergence mechanism
Let us see that learning dynamics in the matching pennies converge to the equilibrium following the mechanism we provided. To analyze the matching pennies in which both the players have two actions, we use the notation of $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$, $q_{j}:=\log x_{j}-\log(1-x_{j})$, and $(y_{1},y_{2})=:(y,1-y)$.
@@ 5.1 Visualization of global convergence mechanism
Fig. 3A shows the dynamics of $q_{1}$ and $q_{2}$. Here, the colors indicate the contour plot for $q_{1}-q_{2}$, showing that $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$ monotonically increases with time and thus Thm. 2 holds. Furthermore, we also see that X’s strategy reaches the region of $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ after a sufficiently long time passes. In the region, $\bm{X}^{\rm T}\bm{U}$ is positive definite, and thus Thm. 1 is applicable after sufficiently long time passes.
Next, Fig. 3B plots the global behavior of the learning dynamics, which is described by the three parameters of $x_{1}$, $x_{2}$, and $y$. The gray line shows the region that corresponds to the Nash equilibrium, i.e., $\bm{x}^{{\rm st}}(\bm{X},\bm{y})=\bm{y}=(1/2,1/2)$. Furthermore, the colored lines show example trajectories of the learning dynamics. The blue part of the line shows that $D(\bm{X},\bm{y})$ increases at the beginning of the learning dynamics. This part is in the region of $x_{1}<x_{2}$, following Thm. 1. After that, the red part shows that $D(\bm{X},\bm{y})$ decreases, and the learning dynamics converge to the equilibrium. This part is in the region of $x_{1}>x_{2}$, following Thm. 1.
5.2 Global convergence in more advanced zero-sum games
The global convergence is observed in more advanced zero-sum games than the matching pennies. Fig. 4 shows three representative examples. In all the examples, we basically assume that there is no dominant (pure-)strategy, where X’s $i=\sigma(j)$-th action is advantageous for Y’s $j$-th action, and vise versa, for the permutation $\sigma$ (i.e., $\sigma(1)=2$, $\sigma(2)=3$, $\sigma(3)=4$, and $\sigma(4)=1$). In other words, the payoff matrices satisfy $u_{i\sigma(i)}=2$ and $u_{\sigma(j)j}=-2$. Three panels take different values in the other elements of their payoff matrix.
@@ 5.2 Global convergence in more advanced zero-sum games
Figure 4: Global convergence in three advanced games. In all the games, we consider the rock-paper-scissors-like game extended to the four-action case, where the second, third, fourth, and first actions win the other’s first, second, third, and fourth actions, respectively. The winner receives the payoff of $2$ (the orange blocks in the matrices for the winning of X), while the loser sends the payoff of $2$ (the blue blocks). We now introduce three variants for the other blocks in the payoff matrix. A. We set each of the other blocks by random numbers in $[-1,1]$ (the gray blocks). Then, Y’s strategy converges to the unique Nash equilibrium (the red star marker) independent of its initial state (the blue circle markers). B. We set each of the other blocks by 00, where the payoff matrix degenerates. Y’s strategy converges to one of the Nash equilibria (the line consisting of the red star markers) depending on its initial state. C. Only the block for the interaction between an action is set to $-1$, and the others are 00. If so, X’s strategy converges to the unique Nash equilibrium (the orange star markers) independent of its initial state (the green circle markers). Instead, Y’s strategies do not converge.

### v2

@@ 3.4. Family of Lyapunov Functions
This means that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ for all $\bm{\delta}$. If we substitute $\bm{\delta}=\bm{x}-\bm{x}^{*}$ for some $\bm{x}\in\Delta^{m_{\sf X}-1}$, we obtain
$H^{\dagger}(\bm{y};\bm{x}-\bm{x}^{*})=|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|^{2}.$ (17)
For any $\bm{y}\neq\bm{y}^{*}\Leftrightarrow{\rm d}\bm{y}\neq\bm{0}$, there is $\bm{x}$ such that $u^{{\rm st}}(\bm{x},\bm{y})-u^{*}\neq 0$. Thus, ${\rm d}\bm{y}=\bm{0}$ is equivalent to $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$. ∎
Let us interpret the function of $H(\bm{X};\bm{\delta})$. First, $H(\bm{X};\bm{\delta})$ is the quadratic form of matrix $\bm{U}\log\bm{X}^{\rm T}$ for $\bm{\delta}$. We now focus on the meaning of $\bm{U}\log\bm{X}^{\rm T}$. The $i^{\prime}$, $i$ element of $\bm{U}\log\bm{X}^{\rm T}$ is given by $\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i}$, in which we denoted vector $\bm{x}_{i}:=(x_{i|j})_{j}$. This is the inner product of the payoff under X taking $i^{\prime}$-th action ($\bm{u}_{i^{\prime}}$) and the logarithmic strategy under X taking $i$-th action ($\log\bm{x}_{i}$). Thus, $\bm{U}\log\bm{X}^{\rm T}$ shows a correspondence matrix between X’s strategy and its payoff matrix. Since $H(\bm{X};\bm{\delta})$ is the quadratic form of the correspondence matrix, it tends to be large when the diagonal elements of the correspondence matrix are large. The diagonal elements, i.e., $\bm{u}_{i}\cdot\log\bm{x}_{i}$, indicate how X exploits Y for each $i$-th action. Thus, $H(\bm{X};\bm{\delta})$ means the degree of exploitation of Y by X.
@@
Figure 3. (A) Trajectories of $q_{1}$ and $q_{2}$. The rainbow contour plot indicates the value of $q_{1}-q_{2}$. All the trajectories monotonically increase $q_{1}-q_{2}$ with time and converge in the area of $q_{1}>q_{2}$ in their final states. (B) Trajectories of the learning dynamics. The black broken line corresponds to the region of Nash equilibria, $\bm{x}^{{\rm st}}=\bm{y}=(1/2,1/2)$. Each colored line shows a trajectory of the learning dynamics. First, the circle markers show the initial states. Following the blue lines, the trajectories diverge from the Nash equilibria ($D(\bm{X},\bm{y})$ increases with time). However, the trajectories stop to diverge and switch to converge to the Nash equilibria ($D(\bm{X},\bm{y})$ decreases), following the red lines. The star markers are the final states and correspond to one of the Nash equilibria.
3.5. Global Behavior by Two Quantities
So far, the two theorems (Thms. 2 and 3) explain how the two quantities vary with time. Let us understand the global behavior of the learning dynamics by interpreting the two quantities.
$D$ explains increasing/decreasing of distance:
First, recall that $D(\bm{X},\bm{y})$ means the distance from the Nash equilibrium. Thm. 2 shows that the distance becomes smaller when $\bm{X}^{\rm T}\bm{U}$ is positive definite, whereas larger when $\bm{X}^{\rm T}\bm{U}$ is negative definite. Here, we focus on $\bm{X}^{\rm T}\bm{U}$ which determines whether the distance becomes smaller or larger. The $j$, $j^{\prime}$ element of $\bm{X}^{\rm T}\bm{U}$ is given by $\bm{x}_{j}\cdot\bm{u}_{j^{\prime}}$, in words, the correspondence between X’s strategy for $j$-th action ($\bm{x}_{j}$) and its payoff for $j^{\prime}$-th action ($\bm{u}_{j^{\prime}}$). The eigenvalues of a matrix are roughly determined by the diagonal elements $\bm{x}_{j}\cdot\bm{u}_{j}$. This diagonal element is larger when $x_{i|j}$ takes a larger value for larger $u_{ij}$, meaning that X exploits Y’s payoff more. As the simplest example, $\bm{X}^{\rm T}\bm{U}$ is positive definite when X takes $x_{i|j}=1$ for $i=\hat{i}$ such that $\hat{i}=\arg\max_{i}u_{ij}$, while $x_{i|j}=0$ for $i\neq\hat{i}$. To summarize, Thm. 2 captures the global behavior where if X exploits Y, their strategies converge to the Nash equilibrium; otherwise, they diverge.
$H$ explains the monotonic increase of exploitability:
Next, recall that $\bm{U}\log\bm{X}^{\rm T}$ in $H(\bm{X};\bm{\delta})$ indicates a correspondence matrix between X’s strategy and its payoff matrix. Thus, Thm. 3 explains that unless Y takes the equilibrium strategy, this correspondence continues to increase with time. We also pay attention to the change in the degree of correspondence, i.e., $H^{\dagger}(\bm{y};\bm{\delta})$. By substituting $\bm{\delta}=\bm{x}-\bm{x}^{*}$ for some $\bm{x}\in\Delta^{m_{\sf X}-1}$, we rewrite it as $H^{\dagger}(\bm{y};\bm{x}-\bm{x}^{*})=|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|^{2}$. Here, $|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|$ indicates the difference in payoff from the equilibrium, i.e., the exploitation by X to Y. Therefore, the correspondence between X’s strategy and its payoff matrix becomes larger according to the exploitability.
Remark:
We have interpreted both $\bm{X}^{\rm T}\bm{U}$ and $\bm{U}\log\bm{X}^{\rm T}$ as the degree of correspondence between X’s strategy and its payoff matrix. The interpretation is qualitatively true, but we remark that there are several quantitative differences between $\bm{X}^{\rm T}\bm{U}$ and $\bm{U}\log\bm{X}^{\rm T}$, such as the order of multiplication and the existence of logarithm.
4. Application: Global Convergence
By combining the global behaviors obtained from Thms. 2 and  3, we expect that the global convergence to the Nash equilibrium occurs regardless of X’s and Y’s initial strategies. Thm. 3 shows that as long as Y’s strategy is out of equilibrium, the correspondence between X’s strategy and its payoff matrix continues to be stronger. Afterward, Thm. 2 shows that if the correspondence is sufficiently strong, Y’s strategy is induced to the Nash equilibrium. In the following, our theory and experiment support that such global convergence occurs.
4.1. Example: Matching Pennies
Let us define the matching pennies game (see Fig. 2 for the illustration of its payoff matrix). This game considers the action numbers of $m_{\sf X}=m_{\sf Y}=2$ and the payoff matrix of $\bm{U}=((u_{11},u_{12}),(u_{21},u_{22}))=((+1,-1),(-1,+1))$. The Nash equilibrium of this game is only $\bm{x}^{*}=\bm{y}^{*}=(1/2,1/2)$. This game is the simplest example of a game equipped with a full-support Nash equilibrium. In addition, it has been known that the replicator dynamics in games without memories cycle around the Nash equilibrium and cannot reach the equilibrium. Nevertheless, by considering the memory asymmetry (Eqs. (5) and (6)), Y’s strategy succeeds in the convergence to the equilibrium, as shown in the following corollary (see Appendix A.3 for its full proof). For convenience, we use a special notation available in two-action games; $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$, $q_{j}:=\log x_{j}-\log(1-x_{j})$, and $(y_{1},y_{2})=:(y,1-y)$.
Corollary 0 (Global convergence in matching pennies).
In the matching pennies game $\bm{U}=((+1,-1),(-1,+1))$, Y’s strategy $\bm{y}$ converges to the equilibrium $\bm{y}^{*}$, independent of both the players’ initial strategies.
Proof Sketch. First, note that $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$. By the direct calculation, we prove $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$. Thm. 3 shows that as long as $\bm{y}=\bm{y}^{*}$, $H(\bm{X};\bm{\delta})$ continues to increase. Thus, after a sufficiently long time, $H(\bm{X};\bm{\delta})>0\Leftrightarrow q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ continue to hold. We can also prove that $x_{1}>x_{2}$ is equivalent to the positive definiteness of $\bm{X}^{\rm T}\bm{U}$. Thm. 2 shows that under positive definite $\bm{X}^{\rm T}\bm{U}$, $\bm{y}$ asymptotically converges to $\bm{y}^{*}$, its equilibrium strategy. ∎
@@ 4.1. Example: Matching Pennies
Our experiments visualize the mechanism of the global convergence based on Thm. 2 and 3. Fig. 3A shows the dynamics of $q_{1}$ and $q_{2}$. Here, the colors indicate the contour plot for $q_{1}-q_{2}$, showing that $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$ monotonically increases with time and thus Thm. 3 holds. Furthermore, we also see that X’s strategy reaches the region of $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ after a sufficiently long time passes. In the region, $\bm{X}^{\rm T}\bm{U}$ is positive definite, and thus Thm. 2 is applicable after sufficiently long time passes.
@@
Figure 4. Global convergence in the coupled matching pennies games, where the second, third, fourth, and first actions win the other’s first, second, third, and fourth actions, respectively. The winner receives the payoff of $2$ (the orange blocks in the matrices for the winning of X), while the loser sends the payoff of $2$ (the blue blocks). We now introduce three variants for the other blocks in the payoff matrix. (A) The case of interior equilibrium. We set each of the other blocks by random numbers in $[-1,1]$ (the gray blocks). Then, Y’s strategy converges to the unique Nash equilibrium (the red star marker) independent of its initial state (the blue circle markers). (B) The case of continuous equilibrium. We set each of the other blocks by 00, where the payoff matrix degenerates. Y’s strategy converges to one of the Nash equilibria (the line consisting of the red star markers) depending on its initial state. (C) The case of boundary equilibrium. Only the block for the interaction between an action is set to $-1$, and the others are 00. If so, X’s strategy converges to the unique Nash equilibrium (the orange star markers) independent of its initial state (the green circle markers). Instead, Y’s strategies do not converge.
Next, Fig. 3B plots the global behavior of the learning dynamics, which is described by the three parameters of $x_{1}$, $x_{2}$, and $y$. The gray line shows the region that corresponds to the Nash equilibrium, i.e., $\bm{x}^{{\rm st}}(\bm{X},\bm{y})=\bm{y}=(1/2,1/2)$. Furthermore, the colored lines show example trajectories of the learning dynamics. The blue part of the line shows that $D(\bm{X},\bm{y})$ increases at the beginning of the learning dynamics. This part is in the region of $x_{1}<x_{2}$, following Thm. 2. After that, the red part shows that $D(\bm{X},\bm{y})$ decreases, and the learning dynamics converge to the equilibrium. This part is in the region of $x_{1}>x_{2}$, following Thm. 2.
4.2. Example: Coupled Matching Pennies
We also observe the global convergence in other zero-sum games beyond the matching pennies game. Fig. 4 considers three examples of “coupled” matching pennies games, where a pair of matching pennies games are coupled with some interaction. The game considers the action numbers of $m_{\sf X}=m_{\sf Y}=4$, and some elements of the payoff matrix is fixed as $u_{ij}=+2$ for $j=\sigma(i)$ and $u_{ij}=-2$ for $i=\sigma(j)$. Here, we used the permutation function $\sigma$ as $\sigma(1)=2$, $\sigma(2)=3$, $\sigma(3)=4$, and $\sigma(4)=1$. If we consider only X’s odd actions and Y’s even actions or its reverse, the payoff matrix corresponds to the matching pennies game. Interestingly, there are various types of Nash equilibrium depending on the interaction between these matching pennies games, i.e., the other elements of the payoff matrix, $u_{ij}$ for neither $j=\sigma(i)$ nor $i=\sigma(j)$. Indeed, Fig. 4 shows three cases where Nash equilibrium exists (A) in the interior, (B) continuously, and (C) on the boundary.
We remark that the global behavior obtained from Thm. 2 and 3 are available even though the trajectories of the learning dynamics look complicated (see Fig. 4). Furthermore, Panels A and B show that Y’s strategy converges to the equilibrium. The only exception is Panel C, but X’s strategy converges to the equilibrium instead of Y’s. In the following, we explain in detail this convergence for each panel.
Interior equilibrium: